5=11+2x-2
2x = 5-11+2=-4
x = -2

Found 2 solutions by don vito, vbeulah:
Answer by don vito(18)

5=11+2x-2
+2
7=11+2x
-11
-4=2x
-----
2
x=-2
there's your friggin' answer. now, leave me alone, po-po!

Answer by vbeulah(1)